# Probing quark-lepton correlation in GUTs with high-precision neutrino measurements

### v1

@@ Appendix A Details of fitting procedure
    We begin with 25000 random points whose parameters $x_{i}^{0}$ (ii runs through free parameters of every models) are within their defined region shown in Tab. 3, from which we obtain Dirac mass matrices, and the light neutrino mass matrix are obtained through type-I dominant seesaw. The initial value of $\chi^{2}$ is usually very large, reaching the order of $10^{8}$.
@@
    We further minimize $\chi^{2}$ by varying parameters of initial points $x_{i}^{0}$ within the range of $x_{i}^{0}(1\pm\varepsilon)$. Here, $\varepsilon$ describes the degree to which the size of the interval parameters $x_{i}^{0}$ can vary in. Furthermore, $\varepsilon$ can be adaptively adjusted by code within the range of $(10^{-4},0.03)$ to achieve smaller $\chi^{2}$. Then, we find a series of points with smaller $\chi^{2}$, whose parameters are denoted as $x_{i}^{1}$.
@@
    We repeat step 2 by replacing $x_{i}^{k+1}$ with $x_{i}^{k}$ and minimize $\chi^{2}$. We run the iteration number $k$ up to $2\times 10^{6}$. Finally, we select points with $\chi^{2}/n_{\rm obs}\leq 10$.
Our scan is performed on the HIAS cluster with 3000 CPU cores available. Model M1 with normal ordering and $\theta_{23}$ in 1st octant (M1N1) or M1 with normal ordering and $\theta_{23}$ in 2nd octant (M1N2) complete their scan using 2048 CPU cores in 16 hours and get final $\chi^{2}/n_{\rm obs}=3.87,3.10$. Model M2 with normal ordering and $\theta_{23}$ in 1st octant (M2N1) or M2 with normal ordering and $\theta_{23}$ in 2nd octant (M2N2) complete their scan using 896 CPU cores in 12 hours and get final $\chi^{2}/n_{\rm obs}=1.34,2.65$. Model M3 with normal ordering and $\theta_{23}$ in 1st octant (M3N1) or M3 with normal ordering and $\theta_{23}$ in 2nd octant (M3N2) complete their scan using 1000 CPU cores in 16 hours and get final $\chi^{2}/n_{\rm obs}=0.27,0.04$. Model M1 with inverted ordering and $\theta_{23}$ in 1st octant (M1I1) or M1 with inverted ordering and $\theta_{23}$ in 2nd octant (M1I2) complete their scan using 768 CPU cores in 9 hours and get final $\chi^{2}/n_{\rm obs}=280.18,280.05$. Model M2 with inverted ordering and $\theta_{23}$ in 1st octant (M2I1) or M2 with inverted ordering and $\theta_{23}$ in 2nd octant (M2I2) complete their scan using 768 CPU cores in 9 hours and get final $\chi^{2}/n_{\rm obs}=34.47,44.98$. Model M3 with inverted ordering and $\theta_{23}$ in 1st octant (M3I1) or M3 with inverted ordering and $\theta_{23}$ in 2nd octant (M3I2) complete their scan using 768 CPU cores in 9 hours and get final $\chi^{2}/n_{\rm obs}=19.14,19.04$.
Table 3: Parameters counting for 3 different models.
Label Free parameters
$\begin{array}[]{c}\text{M1}\\ (n_{\rm para}=19)\end{array}$ $\begin{array}[]{c}\{r_{1},r_{2}\}\in(-10,10),\ m_{0}\in(0,1)\ {\rm eV},\ \omega\in[0,2\pi),\\ D:\{D_{11},D_{22},D_{33}\}\in(0,1),\\ S:\{S_{11}e^{is_{1}},S_{12}e^{is_{2}},S_{13}e^{is_{3}},S_{22}e^{is_{4}},S_{23}e^{is_{5}},S_{33}e^{is_{6}}\}\in(-10,10)e^{i[0,2\pi)}\end{array}$
$\begin{array}[]{c}\text{M2}\\ (n_{\rm para}=19)\end{array}$ $\begin{array}[]{c}\{r_{1},r_{4}\}\in(-10,10),~m_{0}\in(0,1)\ {\rm eV},~\{\alpha,\lambda,\gamma,\omega\}\in[0,2\pi),\\ D:\{D_{11},D_{22},D_{33}\}\in(0,1),\\ S:\{S_{11},S_{12},S_{13},S_{22},S_{23},S_{33}\}\in(-10,10),\\ A:\{A_{12},A_{13},A_{23}\}\in(-10,10)\end{array}$
$\begin{array}[]{c}\text{M3}\\ (n_{\rm para}=18)\end{array}$ $\begin{array}[]{c}\{r_{1},r_{2},r_{3},r_{4},r_{5}\}\in(-10,10),~m_{0}\in(0,1)\ {\rm eV},\\ D:\{D_{11},D_{22},D_{33}\}\in(0,1),\\ S:\{S_{11},S_{12},S_{13},S_{22},S_{23},S_{33}\}\in(-10,10),\\ A:\{A_{12},A_{13},A_{23}\}\in(-10,10)\end{array}$
@@ Appendix B Benchmark studies
We show a few benchmark points in M1, M2 and M3 extracted from the scan in Tab. 4.
@@
Table 4: Benchmark fit values and predictions in 3 models. Each benchmark is named, e.g., BM1N1: Benchmark of M1 with Normal ordering and 1st octant of $\theta_{23}$.
Physical Quantity M1 M2 M3
BM1N1 BM1N2 BM2N1 BM2N2 BM3N1 BM3N2
$y_{u}/10^{-6}$ 4.1841 4.08227 2.81639 7.43413 2.57195 2.51847
$y_{c}/10^{-3}$ 1.40605 1.37123 1.36956 1.38067 1.35808 1.36456
$y_{t}$ 0.428387 0.42744 0.428109 0.427893 0.428165 0.427999
$y_{d}/10^{-6}$ 1.90741 2.33024 3.81711 6.40586 7.63478 6.50436
$y_{s}/10^{-4}$ 1.20008 1.35263 1.17849 1.35147 1.22573 1.22668
$y_{b}/10^{-2}$ 0.56536 0.572574 0.569053 0.569561 0.569812 0.571195
$y_{e}/10^{-6}$ 2.70615 2.70217 2.70694 2.70512 2.7012 2.70299
$y_{\mu}/10^{-4}$ 5.67557 5.64767 5.72306 5.68859 5.70761 5.70903
$y_{\tau}/10^{-2}$ 0.979853 0.971086 0.968229 0.969382 0.968972 0.968018
$\theta_{12}^{q}$ 0.227388 0.227475 0.226987 0.227082 0.22744 0.227423
$\theta_{23}^{q}/10^{-2}$ 4.78722 4.86052 4.8502 4.85534 4.86112 4.86603
$\theta_{13}^{q}/10^{-3}$ 4.44527 4.36218 4.15777 4.03305 4.19948 4.19105
$\delta^{q}$ 1.10098 1.20818 1.17491 1.15692 1.25108 1.23699
$\Delta m_{21}^{2}/10^{-5}~\mathrm{[eV]}$ 7.49642 7.51729 7.48813 7.3983 7.4778 7.49186
$\Delta m_{31}^{2}/10^{-3}~\mathrm{[eV]}$ 2.51096 2.53253 2.51289 2.53819 2.51446 2.53476
$\mathrm{sin}^{2}\theta_{12}^{l}$ 0.306389 0.315291 0.299572 0.28436 0.306322 0.307802
$\mathrm{sin}^{2}\theta_{23}^{l}$ 0.469948 0.551379 0.444945 0.561499 0.484914 0.566208
$\mathrm{sin}^{2}\theta_{13}^{l}$ 0.021243 0.0210321 0.0223411 0.0223394 0.0219516 0.0218592
$\chi^{2}/(n_{\mathrm{obs}}=18)$ 3.87235 3.09649 1.33774 2.64746 0.267685 0.0398901
$m_{1}~\mathrm{[eV]}$ 0.00428924 0.00580787 0.0000445112 0.000277348 0.000786665 0.000344881
$m_{\beta}~\mathrm{[eV]}$ 0.0985221 0.102458 0.0939997 0.09378 0.0940432 0.0940024
$m_{\beta\beta}~\mathrm{[eV]}$ 0.0672365 0.0872486 0.0604685 0.0602027 0.0644098 0.0625105
$\delta~\mathrm{[^{\circ}]}$ 285.682 208.427 350.862 310.236 346.342 339.851
$M_{1}~\mathrm{[GeV]}$ $1.15\times 10^{9}$ $1.43\times 10^{9}$ 60633.6 $2.56\times 10^{6}$ $6.56\times 10^{9}$ $1.26\times 10^{10}$
$M_{2}~\mathrm{[GeV]}$ $1.74\times 10^{11}$ $1.72\times 10^{11}$ $4.28\times 10^{12}$ $7.44\times 10^{11}$ $1.37\times 10^{11}$ $2.28\times 10^{11}$
$M_{3}~\mathrm{[GeV]}$ $3.11\times 10^{12}$ $3.05\times 10^{12}$ $1.29\times 10^{15}$ $2.15\times 10^{14}$ $6.26\times 10^{12}$ $5.52\times 10^{12}$

### v2

@@ Appendix A Details of fitting procedure
    Firstly, we begin with 15000 randomly generated points whose parameters $x_{i}^{0}$ (where i runs over all free parameters of each model) are sampled within their defined region shown in Tab. III. From these parameters, we obtain light neutrino mass matrix via seesaw formula and Dirac mass matrices. The initial value of $\chi^{2}$ is usually very large, of order $10^{8}$.
@@
    Then, we minimize $\chi^{2}$ using the differential evolution (DE) algorithm from the SciPy library. The DE algorithm optimizes all points simultaneously. The configuration uses 15,000 independent chains running in parallel via MPI, with each chain maintaining a population of 15 points. The best1bin mutation strategy ($F=0.7$) is applied for up to 50000 generations.
@@
    Finally, we truncate points with $\chi^{2}/n_{\rm obs}<10$ for each model.
Our scan is performed on the HIAS cluster with 3000 CPU cores available. Code and derived points of all three models are released in GitHub. We further show these points in Fig. 3 in the normal ordering and Fig. 4 in the inverted ordering.
Figure 3: Predictions of models M1, M2 and M3 for $\chi^{2}/n_{\rm obs}<10$ in fitting data N1 and N2 in the normal ordering.
Figure 4: Predictions of models M2 and M3 for $\chi^{2}/n_{\rm obs}<10$ in fitting data I1 and I2 in the inverted ordering. M1 is not shown since no viable points found in the inverted ordering.
In the normal ordering of light neutrino masses, using model M1 to fit data N1 (M1N1) or M1 to fit N2 (M1N2) complete their scan using 1500 CPU cores in 28 hours and minimize the $\chi^{2}$ value to $\chi_{\rm min}^{2}/n_{\rm obs}=1.54,1.71$. M2N1 and M2N2 complete scans with 1500 CPU cores in 13 hours and get $\chi_{\rm min}^{2}/n_{\rm obs}=3.34\times 10^{-6},7.60\times 10^{-6}$. M3N1 and M3N2 complete scans using 3000 CPU cores in 27 hours and get final $\chi_{\rm min}^{2}/n_{\rm obs}=3.41\times 10^{-7},3.19\times 10^{-7}$.
In the inverted ordering, we fit I1 and I2 in model M1 with 1500 CPU cores in 24 hours, and obtain $\chi_{\rm min}^{2}/n_{\mathrm{obs}}=147.41$ and $147.64$, respectively, thus, no viable point for model M1 found in the inverted ordering. In model M2, M2I1 and M2I2 complete scans with 1500 CPU cores in 12 hours and obtain $\chi_{\rm min}^{2}/n_{\mathrm{obs}}=0.205$ and $0.00721$, respectively. M3I1 and M3I2 complete scans using 1500 CPU cores in 26 hours and obtain $\chi_{\rm min}^{2}/n_{\mathrm{obs}}=0.0680$ and $0.598$, respectively.
We comment on the prediction of light neutrino masses which are not discussed in the maintext. In the normal ordering, the lightest neutrino mass $m_{1}$ is predicted around the region $[0.002,0.01]$ eV in M1, $(0,0.015]$ eV in M2 and mainly distributed in $(0,0.02]$ eV in M3. $m_{\beta\beta}$ in M1 is predicted in a narrow regime $[0.004,0.01]$ eV, which can be tested in the future $0\nu\beta\beta$ experiments. M2 prefers a smaller value, mainly in the region $[0.003,0.006]$ eV, below the experimental sensitivity. The predicted region in M3 overlaps with both models. In the inverted ordering, the lightest neutrino mass $m_{3}$ is smaller than 0.02 eV and $0\nu\beta\beta$ mass $m_{\beta\beta}\gtrsim 0.05$ eV in our limited data points.
@@ Appendix B Benchmark studies
We show benchmark points of M1, M2 and M3 in Tab. 4 and 5 with four versions of data N1, N2, I1 and N2 all taken into account. In models M2 and M3, as diverse mass hierarchies for RHNs are predicted, we show a few distinguishable benchmarks in these models. For example, very large hierarchy $(M_{1},M_{2},M_{3})=(9.95\cdot 10^{4},2.14\cdot 10^{12},7.51\cdot 10^{14})$ GeV is predicted in M2N1B1, which is consistent with the ansatze assumed in Babu:2024ahk , relatively smaller hierarchy $(M_{1},M_{2},M_{3})=(7.36\cdot 10^{6},6.96\cdot 10^{10},2.29\cdot 10^{13})$ GeV is also found in M2N1B2 in our scan.
Table 4: Benchmark fit values and predictions in 3 models in the normal ordering. Each benchmark is named, e.g., M1N1B1: benchmark 1 of M1 in fitting data N1. $\Delta m^{2}_{3l}$ = $\Delta m^{2}_{31}>0$ and $m_{l}=m_{1}$ denote the lightest neutrino mass. We distinguish different benchmarks, e.g., M1N1B1 vs M1N1B2, due to the different hierarchies of RHN masses are predicted when fitting the same data in the same model.
Physical M1 M2 M3
Quantity M1N1B1 M1N2B1 M2N1B1 M2N1B2 M2N2B1 M2N2B2 M3N1B1 M3N1B2 M3N2B1 M3N2B2
$y_{u}/10^{-6}$ 2.65 2.75 2.55 3.07 2.53 4.85 0.637 2.54 0.744 2.54
$y_{c}/10^{-3}$ 1.40 1.39 1.37 1.35 1.37 1.37 1.37 1.37 1.37 1.37
$y_{t}$ 0.429 0.427 0.428 0.429 0.428 0.430 0.428 0.428 0.428 0.428
$y_{d}/10^{-6}$ 3.87 3.13 6.56 7.29 6.56 7.14 5.28 6.56 5.28 6.56
$y_{s}/10^{-4}$ 1.37 1.32 1.24 1.17 1.24 1.11 1.24 1.24 1.25 1.24
$y_{b}/10^{-2}$ 0.571 0.571 0.570 0.570 0.570 0.568 0.569 0.570 0.569 0.570
$y_{e}/10^{-6}$ 2.69604 2.70309 2.70346 2.70190 2.70339 2.70629 2.70318 2.70346 2.70345 2.70341
$y_{\mu}/10^{-4}$ 5.65981 5.68560 5.70708 5.72116 5.70713 5.70725 5.71839 5.70706 5.71432 5.70698
$y_{\tau}/10^{-2}$ 0.97358 0.97054 0.97021 0.96623 0.97020 0.97211 0.97154 0.97020 0.97124 0.97020
$\theta_{12}^{q}$ 0.22750 0.22749 0.22739 0.22760 0.22739 0.22721 0.22745 0.22739 0.22745 0.22739
$\theta_{23}^{q}/10^{-2}$ 4.904 4.872 4.858 4.876 4.858 4.835 4.858 4.858 4.856 4.858
$\theta_{13}^{q}/10^{-3}$ 4.252 4.215 4.202 4.201 4.202 4.237 4.202 4.202 4.202 4.202
$\delta^{q}$ 1.269 1.251 1.207 1.253 1.207 1.227 1.201 1.207 1.203 1.207
$\Delta m_{21}^{2}/10^{-5}\,\mathrm{eV}^{2}$ 7.53 7.50 7.50 7.50 7.50 7.48 7.50 7.50 7.50 7.50
$\Delta m_{3\ell}^{2}/10^{-3}\,\mathrm{eV}^{2}$ 2.515 2.535 2.513 2.512 2.534 2.539 2.513 2.513 2.534 2.534
$\mathrm{sin}^{2}\theta_{12}^{l}$ 0.313 0.309 0.309 0.303 0.309 0.308 0.309 0.309 0.309 0.309
$\mathrm{sin}^{2}\theta_{23}^{l}$ 0.456 0.561 0.470 0.460 0.561 0.540 0.471 0.470 0.561 0.561
$\mathrm{sin}^{2}\theta_{13}^{l}$ 0.02152 0.02180 0.02215 0.02235 0.02195 0.02223 0.02217 0.02215 0.02194 0.02195
$\chi^{2}/(n_{\mathrm{obs}}=18)$ 1.54 1.71 $5.88\cdot 10^{-6}$ 0.312 $7.60\cdot 10^{-6}$ 0.958 0.495 $3.41\cdot 10^{-7}$ 0.462 $3.39\cdot 10^{-7}$
$m_{l}/\mathrm{eV}$ $3.34\cdot 10^{-3}$ $5.60\cdot 10^{-3}$ $5.29\cdot 10^{-5}$ $7.81\cdot 10^{-4}$ $5.41\cdot 10^{-5}$ $1.52\cdot 10^{-3}$ $3.82\cdot 10^{-3}$ $5.46\cdot 10^{-4}$ $3.44\cdot 10^{-3}$ $1.03\cdot 10^{-3}$
$m_{\beta}/\mathrm{eV}$ $9.40\cdot 10^{-3}$ $1.05\cdot 10^{-2}$ $85\cdot 10^{-3} 8.8$$8\cdot 10^{-3} 8.8$$85\cdot 10^{-3}$ $9.02\cdot 10^{-3}$ $9.64\cdot 10^{-3}$ $8.87\cdot 10^{-3}$ $9.49\cdot 10^{-3}$ $8.91\cdot 10^{-3}$
$m_{\beta\beta}/\mathrm{eV}$ $4.94\cdot 10^{-3}$ $7.18\cdot 10^{-3}$ $3.70\cdot 10^{-3}$ $4.22\cdot 10^{-3}$ $3.67\cdot 10^{-3}$ $4.76\cdot 10^{-3}$ $6.51\cdot 10^{-3}$ $4.09\cdot 10^{-3}$ $6.17\cdot 10^{-3}$ $4.42\cdot 10^{-3}$
$\delta/\mathrm{{}^{\circ}}$ 38.816 210.46 43.046 11.532 341.29 6.4034 186.19 14.663 189.07 351.61
$M_{1}/\mathrm{GeV}$ $7.48\cdot 10^{9}$ $2.56\cdot 10^{9}$ $9.95\cdot 10^{4}$ $7.36\cdot 10^{6}$ $8.43\cdot 10^{4}$ $7.13\cdot 10^{6}$ $1.01\cdot 10^{3}$ $1.37\cdot 10^{10}$ $9.09\cdot 10^{2}$ $2.71\cdot 10^{10}$
$M_{2}/\mathrm{GeV}$ $4.95\cdot 10^{11}$ $2.14\cdot 10^{11}$ $2.14\cdot 10^{12}$ $6.96\cdot 10^{10}$ $3.64\cdot 10^{12}$ $9.07\cdot 10^{10}$ $6.36\cdot 10^{11}$ $2.57\cdot 10^{11}$ $5.60\cdot 10^{11}$ $5.01\cdot 10^{11}$
$M_{3}/\mathrm{GeV}$ $5.61\cdot 10^{12}$ $2.48\cdot 10^{12}$ $7.51\cdot 10^{14}$ $2.29\cdot 10^{13}$ $1.02\cdot 10^{15}$ $2.97\cdot 10^{13}$ $1.36\cdot 10^{13}$ $9.45\cdot 10^{12}$ $1.19\cdot 10^{13}$ $2.09\cdot 10^{13}$
@@
Table 5: Benchmark fit values and predictions in 2 models in the inverted ordering. Each benchmark is named, e.g., M2I2B1: Benchmark 1 of M2 in fitting data I2. $\Delta m^{2}_{3\ell}$ = $\Delta m^{2}_{32}<0$ and $m_{l}=m_{3}$ denote the lightest neutrino mass.
Physical M2 M3
Quantity M2I1B1 M2I1B2 M2I2B1 M2I2B2 M3I1B1 M3I1B2 M3I2B1 M3I2B2
$y_{u}/10^{-6}$ 2.79 4.03 2.77 3.49 2.56 2.54 2.56 2.54
$y_{c}/10^{-3}$ 1.37 1.44 1.37 1.36 1.38 1.36 1.38 1.36
$y_{t}$ 0.428 0.427 0.428 0.428 0.428 0.428 0.428 0.428
$y_{d}/10^{-6}$ 6.92 6.82 6.60 8.92 5.95 8.05 5.95 7.93
$y_{s}/10^{-4}$ 1.27 0.740 1.24 1.11 1.21 1.10 1.22 1.11
$y_{b}/10^{-2}$ 0.572 0.570 0.570 0.567 0.571 0.570 0.571 0.570
$y_{e}/10^{-6}$ 2.70369 2.70455 2.70332 2.70702 2.70333 2.70326 2.70337 2.70348
$y_{\mu}/10^{-4}$ 5.71365 5.73366 5.70814 5.68800 5.70707 5.71460 5.70701 5.71393
$y_{\tau}/10^{-2}$ 0.96303 0.97624 0.96926 0.98031 0.97013 0.97102 0.97008 0.97048
$\theta_{12}^{q}$ 0.22740 0.22733 0.22739 0.22735 0.22742 0.22741 0.22742 0.22740
$\theta_{23}^{q}/10^{-2}$ 4.864 4.884 4.859 4.860 4.870 4.871 4.870 4.869
$\theta_{13}^{q}/10^{-3}$ 4.140 4.194 4.195 4.036 4.201 4.146 4.200 4.152
$\delta^{q}$ 1.210 1.157 1.206 1.023 1.216 1.217 1.216 1.214
$\Delta m_{21}^{2}/10^{-5}\,\mathrm{eV}^{2}$ 7.49 7.46 7.49 7.28 7.49 7.50 7.49 7.49
$\Delta m_{3\ell}^{2}/10^{-3}\,\mathrm{eV}^{2}$ 2.484-2.484 2.483-2.483 2.510-2.510 2.509-2.509 2.484-2.484 2.484-2.484 2.510-2.510 2.510-2.510
$\mathrm{sin}^{2}\theta_{12}^{l}$ 0.308 0.306 0.308 0.357 0.308 0.308 0.308 0.308
$\mathrm{sin}^{2}\theta_{23}^{l}$ 0.569 0.556 0.565 0.520 0.550 0.567 0.562 0.578
$\mathrm{sin}^{2}\theta_{13}^{l}$ 0.02232 0.02228 0.02224 0.02241 0.02231 0.02237 0.02224 0.02230
$\chi^{2}/(n_{\mathrm{obs}}=18)$ 0.205 4.31 0.00721 3.54 0.0680 0.717 0.0680 0.598
$m_{l}/\mathrm{eV}$ $2.53{\cdot}10^{-4}$ $1.78{\cdot}10^{-2}$ $2.07{\cdot}10^{-4}$ $2.02{\cdot}10^{-2}$ $9.83{\cdot}10^{-3}$ $1.48{\cdot}10^{-2}$ $9.60{\cdot}10^{-3}$ $1.41{\cdot}10^{-2}$
$m_{\beta}/\mathrm{eV}$ $4.88{\cdot}10^{-2}$ $5.19{\cdot}10^{-2}$ $4.90{\cdot}10^{-2}$ $5.31{\cdot}10^{-2}$ $4.97{\cdot}10^{-2}$ $5.10{\cdot}10^{-2}$ $5.00{\cdot}10^{-2}$ $5.10{\cdot}10^{-2}$
$m_{\beta\beta}/\mathrm{eV}$ $4.82{\cdot}10^{-2}$ $5.05{\cdot}10^{-2}$ $4.83{\cdot}10^{-2}$ $5.19{\cdot}10^{-2}$ $4.94{\cdot}10^{-2}$ $5.00{\cdot}10^{-2}$ $4.96{\cdot}10^{-2}$ $5.00{\cdot}10^{-2}$
$\delta/\mathrm{{}^{\circ}}$ 307.27 101.18 326.05 69.966 327.06 347.86 202.42 14.403
$M_{1}/\mathrm{GeV}$ $7.60{\cdot}10^{3}$ $4.73{\cdot}10^{6}$ $5.12{\cdot}10^{3}$ $1.84{\cdot}10^{6}$ $5.90{\cdot}10^{5}$ $7.09{\cdot}10^{9}$ $5.93{\cdot}10^{5}$ $7.14{\cdot}10^{9}$
$M_{2}/\mathrm{GeV}$ $2.15{\cdot}10^{12}$ $1.45{\cdot}10^{10}$ $2.97{\cdot}10^{12}$ $1.67{\cdot}10^{10}$ $7.07{\cdot}10^{10}$ $8.17{\cdot}10^{10}$ $7.14{\cdot}10^{10}$ $8.39{\cdot}10^{10}$
$M_{3}/\mathrm{GeV}$ $6.55{\cdot}10^{14}$ $4.30{\cdot}10^{12}$ $8.61{\cdot}10^{14}$ $5.22{\cdot}10^{12}$ $3.14{\cdot}10^{12}$ $6.45{\cdot}10^{12}$ $3.17{\cdot}10^{12}$ $6.53{\cdot}10^{12}$